# Generic groups and the weak amalgamation property

### v2

@@ 2 Seeking generics
In this section we give examples of group varieties $U$ such that the corresponding $\mathcal{G}_{U}$ contains generic groups. In particular we prove Theorem 2.7 formulated in the introduction. In Section 2.1 we introduce the crucial notion of t-isolation and show how it works for proving WAP.
@@
The following definition concerns cases I - II of Section 1.2.
@@
Let $\bar{c}$ be a subtuple of $\bar{c}_{1}$ and $D(\bar{c}_{1})\in\mathcal{B}_{T}$. We say that $D(\bar{c}_{1})$ is t-isolating (term-isolating) for $\bar{c}$ if the group $\langle\bar{c}\cap\mathsf{Gp}\rangle$ satisfies the same group equalities of the form $w(\bar{c}^{\prime})=1$ with $\bar{c}^{\prime}\subseteq\bar{c}\cap\mathsf{Gp}$ in every member of $[D(\bar{c}_{1})]$.
From now on in this section we assume the circumstances of the basic case, i.e. case I. In particular the structure $M$ has only one sort $\mathsf{Gp}$.
Let $G$ be a subgroup of an enumerated group and $G=\langle\bar{c}\rangle$ for some tuple $\bar{c}\subseteq\omega$. We say that $G$ is t-isolated by $D(\bar{c}_{1})$ if $D(\bar{c}_{1})$ is t-isolating for $\bar{c}$ and $G$ is the corresponding group $\langle\bar{c}\rangle$.
@@
We say that an abstract group $G$ is t-isolated if it has a $\mathsf{Gp}$-copy, say $\langle\bar{c}\rangle$, which (under Case I) is isolated by some diagram $D(\bar{c}_{1})$ for $\bar{c}$ where $\bar{c}_{1}\subseteq\langle\bar{c}\rangle$ in every member of $[D(\bar{c}_{1})]$.
Note that when the condition above holds it can be arranged that $\bar{c}_{1}$ is presented by the same tuple of words on $\bar{c}$ in all members of $[D(\bar{c}_{1})]$.
We say that t-isolated diagrams are dense in $\mathcal{B}_{T}$ if any $D_{0}(\bar{c})\in\mathcal{B}_{T}$ extends to some $D(\bar{c}_{1})$ which is t-isolating for $\bar{c}$ and $\bar{c}_{1}\subseteq\langle\bar{c}\rangle$ in $[D(\bar{c}_{1})]$.
@@
Proof. Necessity. Assuming that $G_{0}$ has a copy, say $\langle\bar{c}_{0}\rangle$, which is t-isolated by some $D_{0}(\bar{c}^{\prime}_{0})$ with $\bar{c}^{\prime}_{0}\subseteq\langle\bar{c}_{0}\rangle$, find an extension $D(\bar{c})\supseteq D_{0}(\bar{c}^{\prime}_{0})$ as in the formulation of WAP. Extend it to some t-isolating $D^{\prime}(\bar{c}^{\prime})$ for $\bar{c}$ with $\bar{c}^{\prime}\subseteq\langle\bar{c}\rangle$. It is easy to see that $D^{\prime}(\bar{c}^{\prime})$ witnesses WAP for $D_{0}(\bar{c}^{\prime}_{0})$. Let $G_{1}$ be the corresponding group $\langle\bar{c}\rangle$. Having $G_{2}=\langle\bar{c}_{2}\rangle$ and $G_{3}=\langle\bar{c}_{3}\rangle$ which contain $G_{1}$ and are t-isolated by some $D^{\prime}_{2}(\bar{c}^{\prime}_{2})$ and $D^{\prime}_{3}(\bar{c}^{\prime}_{3})$ with $\bar{c}^{\prime}_{2}\subseteq\langle\bar{c}_{2}\rangle$ and $\bar{c}^{\prime}_{3}\subseteq\langle\bar{c}_{3}\rangle$ (and where we may assume $\bar{c}\subseteq\bar{c}_{2}\cap\bar{c}_{3}$), we apply WAP for amalgamation $D^{\prime}_{2}(\bar{c}^{\prime}_{2})$ and $D^{\prime}_{3}(\bar{c}^{\prime}_{3})$ over $D_{0}(\bar{c}^{\prime}_{0})$. Assume that $D_{4}(\bar{c}_{4})$ is the corresponding amalgam. Then $\langle\bar{c}_{4}\rangle$ is a group containing $\bar{c}_{2}\cup\bar{c}_{3}$. By t-isolation these elements generate subgroups which are copies of $G_{2}$ and $G_{3}$.
Sufficiency. Let $D(\bar{c})\in\mathcal{B}_{T}$. Extend it to a t-isolating $D_{0}(\bar{c}^{\prime})$ for $\bar{c}$ with $\bar{c}^{\prime}\subseteq\langle\bar{c}\rangle$. Let $G_{0}=\langle\bar{c}\rangle$ be the corresponding t-isolated group. Find $G_{1}=\langle\bar{c}_{1}\rangle$ as in the formulation and the corresponding t-isolating diagram $D_{1}(\bar{c}^{\prime}_{1})$ with $\bar{c}^{\prime}_{1}\subseteq\langle\bar{c}_{1}\rangle$. We may assume that $\bar{c}^{\prime}\subseteq\bar{c}_{1}$. In order to verify amalgamation of extensions of $D_{1}(\bar{c}^{\prime}_{1})$ we firstly note that by density of t-isolating diagrams we may restrict ourselves only to extensions which are t-isolating for some subtuples of their domains. If $D_{2}(\bar{c}^{\prime}_{2})$ and $D_{3}(\bar{c}^{\prime}_{3})$ are such extensions and $G_{2}=\langle\bar{c}_{2}\rangle$ and $G_{3}=\langle\bar{c}_{3}\rangle$ are the corresponding groups with $\bar{c}^{\prime}_{2}\subseteq\langle\bar{c}_{2}\rangle$ and $\bar{c}^{\prime}_{3}\subseteq\langle\bar{c}_{3}\rangle$, then by t-isolation, $G_{1}$ is a subgroup of both $G_{2}$ and $G_{3}$. Find $G_{4}$ which amalgamates $G_{2}$ and $G_{3}$ over $G_{0}$. It obviously belongs to some $[D_{4}(\bar{c}_{4})]$ where $D_{4}(\bar{c}_{4})$ amalgamates $D_{2}(\bar{c}_{2})$ and $D_{3}(\bar{c}_{3})$ over $D_{0}(\bar{c}^{\prime})$. $\Box$
@@
It is worth noting that a slightly weaker form of the density assumption follows from WAP. Indeed, if there is $D(\bar{c})\in\mathcal{B}_{T}$ such that for every $D^{\prime}(\bar{c}^{\prime})\in\mathcal{B}_{T}$ extending $D(\bar{c})$ the diagram $D^{\prime}(\bar{c}^{\prime})$ is not t-isolating for $\bar{c}$ (where $\bar{c}^{\prime}\subseteq\langle\bar{c}\rangle$ is not assumed), then for any such $D^{\prime}(\bar{c}^{\prime})$ two distinct groups can be realized as $\langle\bar{c}\rangle$ in some members of $[D^{\prime}(\bar{c}^{\prime})]$. Thus $\mathcal{B}_{T}$ does not have WAP.
@@
Let us note that any finite group $F=\langle\bar{c}\rangle$ which extends to an enumerated group of $T$, is t-isolated. Indeed, let $\bar{c}^{\prime}$ be an enumeration of $F$. The multiplication table of $F$ can be viewed as the corresponding diagram $D^{\prime}(\bar{c}^{\prime})$ and we also have $\bar{c}^{\prime}\subseteq\langle\bar{c}\rangle$. Using this we can conclude that if $T$ defines a variety of groups and all finitely generated groups satisfying $T$ are residually finite then t-isolated diagrams are dense in $\mathcal{B}_{T}$. Indeed, assuming that $D(\bar{c})\in\mathcal{B}_{T}$ take a finite group generated by $\bar{c}$ and satisfying all relations which appear in $D(\bar{c})$. Let $\bar{c}^{\prime}$ be an enumeration of this group. Then the diagram $D^{\prime}(\bar{c}^{\prime})$ as above works as an appropriate t-isolating extension of $D(\bar{c})$.
We also add here that under the assumptions of Remark 2.5, t-isolated groups satisfying $T$ are exactly finite ones. Indeed, if $G=\langle\bar{c}\rangle$ is infinite then any basic set $[D^{\prime}(\bar{c}^{\prime})]$ containing it also contains a finite group.
@@ 2.2 Abelian groups
Abelian groups form a closed subspace of $\mathcal{G}$. We denote it by $\mathcal{G}_{Abel}$ where $Abel$ denotes the corresponding axiomatization. Under the scenario of case I of Section 1.2 we have $\mathcal{G}_{Abel}=\mathcal{X}_{M,Abel}$
The following theorem is proved in [5]. We give an easy and independent proof of it.
@@
Firstly note that since abelian groups are residually finite Remark 2.5 implies that t-isolated diagrams are dense in $\mathcal{B}_{Abel}$.
Let us verify the property that every t-isolated $G_{0}$ can be extended to a t-isolated $G_{1}$ such that any two t-isolated extensions of $G_{1}$ can be amalgamated over $G_{0}$. As we already mentioned after Remark 2.5 t-isolated abelian groups are exactly finite ones. Now the key point is the fact that the variety $\mathbb{G}_{Abel}$ has the amalgamation property. Indeed, if $A_{0}$ is a subgroup of $A_{1}$ and $\phi$ is an embedding of $A_{0}$ into $A_{2}$, then the group $(A_{1}\oplus A_{2})/\{(x-\phi(x))|x\in A_{0}\}$ is the corresponding amalgamation. $\Box$
2.3 Nilpotent group of prime exponent $p$ with $p>2$
Let $c,p\in\omega\setminus\{0,1\}$ and $p$ be prime $>\mathsf{max}(2,c)$. Nilpotent groups of degree $c$ and of exponent $p$ form a closed subspace of $\mathcal{G}$. We denote it by $\mathcal{G}_{c_{nil}p}$.
@@ 2.3 Nilpotent group of prime exponent $p$ with $p>2$
Proof. Consider $\mathcal{G}_{c_{nil}p}$ as the space $\mathcal{X}_{M,c_{nil}p}$ under the scenario of case I. According to Theorem 1.2 of [7] we have to verify that JEP and WAP hold for $\mathcal{B}_{c_{nil}p}$. Since nilpotent groups of degree $c$ and of exponent $p$ form a variety (we call it $\mathbb{G}_{c_{nil}p}$), JEP is satisfied: apply $\oplus$ for domains of the corresponding diagrams. In order to verify WAP we will apply Proposition 2.3.
Firstly note that since nilpotent groups are residually finite (see Section 5.2 in [4]) t-isolated diagrams are dense in $\mathcal{B}_{c_{nil}p}$. Furthermore, t-isolated groups of the variety $\mathbb{G}_{c_{nil}p}$ are exactly finite ones, see Remark 2.5 and the comment after it. Thus Proposition 2.3 can be applied to finite groups. However, we need some additional algebraic material.
@@
Proof. To satisfy the conclusion of the lemma for $i=1$ take $G_{1}=G_{0}$. We now apply induction. Let $k<c$ and assume that we already have some $H\geq G_{0}$ satisfying the statement for all $i\leq k$. Let $G_{1}$ be a finite extension of $H$ with maximal $\Gamma_{k+1}(G_{1})\cap G_{0}$. Since for all $i\leq k$ we have $\Gamma_{i}(G_{1})\cap G_{0}=\Gamma_{i}(H)\cap G_{0}$, the statement of the conclusion of the lemma holds for $G_{1}$ and all $i\leq k+1$. $\Box$
In [2] A. Baudisch defines expansions of groups from $\mathbb{G}_{c_{nil}p}$ by additional unary predicates $P_{2},\ldots,P_{c}$. In order to present these expansions let $G$ be a group in $\mathbb{G}_{c_{nil}p}$ and let $Q_{1}(G)=P_{1}(G)=G$. When $11<n\leq c+$ 1 let $Q_{n}(G)=\langle\bigcup_{\ell+k=n}[P_{\ell}(G),P_{k}(G)]\rangle$ and $P_{n}(G)$ be a subgroup of $G$ with $Q_{n}(G)\subseteq P_{n}(G)\subseteq Z_{c+1-n}(G)$ and $P_{n}(G)\subseteq P_{n-1}(G)$. Let $\mathbb{G}^{P}_{c,p}$ be the class of all groups in $\mathbb{G}_{c_{nil}p}$ with additional predicates $P_{2}(G),\ldots,P_{c}(G)$ as described above. Let $L^{+}$ be the corresponding language that extends the language of groups. Note that if we extend a group $G\in\mathbb{G}_{c_{nil}p}$ by predicates $Q_{i}(G)=P_{i}(G)=\Gamma_{i}(G)$ for $1<i\leq c$ then we get an $L^{+}$-group from $\mathbb{G}^{P}_{c,p}$. The following statement is the main result of [2] (see Theorem 1.2 and the discussion after it).
The class of finite structures from $\mathbb{G}^{P}_{c,p}$ has hereditary property and the amalgamation property.
@@
Consider $G_{1}$, $G_{2}$ and $G_{3}$ as $L^{+}$-structures defining $P_{i}$ on them by $\Gamma_{i}(G_{1})$, $\Gamma_{i}(G_{2})$ and $\Gamma_{i}(G_{3})$ respectively ($1<i\leq c$). Now for any $i\leq c$ let $P_{i}(G_{0})=\Gamma_{i}(G_{1})\cap G_{0}$. By hereditary property $(G_{0},P_{2}(G_{0}),\ldots,P_{c}(G_{0}))$ is also an $L^{+}$-structure from $\mathbb{G}^{P}_{c,p}$.
Since $G_{1}$ is a subgroup of $G_{2}$ and $G_{3}$ then by Lemma 2.8, for any $i\leq c$, we have that $P_{i}(G_{0})=P_{i}(G_{1})\cap G_{0}=P_{i}(G_{2})\cap G_{0}=P_{i}(G_{3})\cap G_{0}$. In particular $(G_{0},P_{2}(G_{0}),\ldots,P_{c}(G_{0}))$ is an $L^{+}$-substructure of both $(G_{2},P_{2}(G_{2}),\ldots,P_{c}(G_{2}))$ and $(G_{3},P_{2}(G_{3}),\ldots,P_{c}(G_{3}))$. By Theorem 1.2 of [2] there is $L^{+}$-structure $G_{4}$ amalgamating $G_{2}$ and $G_{3}$ over $G_{0}$. Thus the condition of Proposition 2.3 is satisfied. $\Box$

### v3

@@ 2 Seeking generics
In this section, we give examples of group varieties $U$ such that the corresponding $\mathcal{G}_{U}$ contains generic groups. In particular, we prove Theorem 2.9 formulated in the introduction. In Section 2.1 we introduce the crucial notion of t-isolation and show how it works for proving WAP.
@@
Consider a countable atomic structure $M$ in a language $L$. Let $T$ be an expansion of $Th(M)$ in some $L\cup\bar{r}$ where $\bar{r}=(r_{1},\ldots,r_{\iota})$ is a sequence of additional relational/functional symbols. We assume that $T$ satisfies all the assumptions stated in the beginning of Section 1.1. Let $M$ have a distinguished sort $\mathsf{Gp}$. Under Scenarios $I$ or $II$ of Section 1.2, it is assumed that $\mathsf{Gp}$ is a sort for a group. However, in Definition 2.1 we do not take this assumption.
Let $\bar{c}$ be a tuple of the sort $\mathsf{Gp}$ of the structure $M$. For any expansion $(M,{\bf\bar{r}})$ the tuple $\bar{c}$ determines a theory, say $\Xi_{\,\bar{c},{\bf\bar{r}}}$, which consists of all valid conditions $r_{i}(t_{1}(\bar{c}_{1}),\ldots,t_{k}(\bar{c}_{k}))$ (or their negations) and $t_{1}(\bar{c}^{\prime})=(\not=)t_{2}(\bar{c}^{\prime\prime})$, where $\bar{c}_{1},\ldots,\bar{c}_{k}$, $\bar{c}^{\prime}$ and $\bar{c}^{\prime\prime}$ are subtuples of $\bar{c}$ and $t_{1}(\bar{x}_{1}),\ldots,t_{k}(\bar{x}_{k})$, $t_{1}(\bar{x}^{\prime})$ and $t_{2}(\bar{x}^{\prime\prime})$ are terms of $L\cup\bar{r}$. In the following definition we distinguish the case when $\Xi_{\,\bar{c},\bf\bar{r}}$ is uniquely determined by some diagram.
@@
Let $\bar{c}$ be a subtuple of $\bar{d}$ and $D(\bar{d})\in\mathcal{B}_{T}$ (and we do not assume that $\bar{c}$ is of the sort $\mathsf{Gp}$). We say that $D(\bar{d})$ is t-isolating (term-isolating) for $\bar{c}$ if the theory $\Xi_{\,\bar{c}\cap\mathsf{Gp},{\bf\bar{r}}}$ is the same in every member $(M,{\bf\bar{r}})\in[D(\bar{d})]$.
In this definition $\bar{c}\cap\mathsf{Gp}$ denotes the subtuple of $\bar{c}$ of all entries which belong to $\mathsf{Gp}$.
When $(M,{\bf\bar{r}})$ is an expansion of $M$ as in the definition, we may consider $\mathsf{Gp}$ as a structure with respect to those relations/functions of $L\cup\bar{r}$ which do not depend on elements of other sorts. Then any subset of $\mathsf{Gp}$ closed under operations of $\mathsf{Gp}$ of this kind is viewed as a substructure. It is worth noting here, that the theory $\Xi_{\,\bar{c},\bf\bar{r}}$ defined before the definition is the same with the corresponding theory defined in this substructure.
Definition 2.1 will be especially important in cases I - II of Section 1.2. In that cases, we may assume that the theory $\Xi_{\,\bar{c}\cap\mathsf{Gp},{\bf\bar{r}}}$ consists of all valid conditions $w(\bar{c}^{\prime})=(\not=)1$ where $\bar{c}^{\prime}$ is a subtuple of $\bar{c}\cap\mathsf{Gp}$ and $w(\bar{x}^{\prime})$ is a group word. Then it is worth noting that Definition 2.1 implies that the tuple $\bar{c}$ generates the same marked group $\langle\bar{c}\cap\mathsf{Gp}\rangle$ in all members $(M,{\bf\bar{r}})\in[D(\bar{d})]$ (as a subgroup of the group defined on $\mathsf{Gp}$).
From now on in this section we assume the circumstances of the basic case, i.e. case I. In particular, the structure $M$ has only one sort $\mathsf{Gp}$. Then, $\Xi_{\,\bar{c}\cap\mathsf{Gp},{\bf\bar{r}}}=\Xi_{\,\bar{c},{\bf\bar{r}}}$ in the definition.
Let $G$ be a subgroup of an enumerated group and $G=\langle\bar{c}\rangle$ for some tuple $\bar{c}\subseteq\omega$. We say that $G$ is t-isolated by $D(\bar{d})$, if $D(\bar{d})$ is t-isolating for $\bar{c}$, and $G$ is the corresponding subgroup $\langle\bar{c}\rangle$.
Example 2.2
Let $F=\langle\bar{c}\rangle$ be a finite group which extends to an enumerated group of $T$. Let $\bar{d}$ be an enumeration of $F$. The multiplication table of $F$ together with inequalities between distinct elements, can be viewed as a diagram $D(\bar{d})$ which is t-isolating for $\bar{c}$.
Now it is easy to see that every finite group which can be embedded into an enumerated group of $T$ is a t-isolated group in the sense of the following definition.
@@
We say that an abstract group $G$ is t-isolated with respect to $T$ if there is an enumerated group, say $G_{1}$, which is defined on $\mathsf{Gp}$ (under Case I) and there is a tuple $\bar{c}\subset\omega$ such that the subgroup $\langle\bar{c}\rangle\leq G_{1}$, is isolated by some diagram $D(\bar{c}_{1})$ for $\bar{c}$, where $\bar{c}_{1}\subseteq\langle\bar{c}\rangle$ in every member of $D(\bar{c}_{1})]$.
Note that when the condition above holds, then extending the diagram if necessary, it can be arranged that $\bar{c}_{1}$ is presented by the same tuple of words on $\bar{c}$ in all members of $D(\bar{c}_{1})]$, and, furthermore, $\bar{c}$ is a subtuple of $\bar{c}_{1}$. Another easy observation is the fact that the set of all enumerated groups from $\mathcal{X}_{M,T}$ which contain a t-isolated $G$ (with a fixed set of generators $\bar{c}$), form a clopen subset of $\mathcal{X}_{M,T}$.
Definition 2.4
We say that t-isolated diagrams are dense in $\mathcal{B}_{T}$ if any $D_{0}(\bar{c})\in\mathcal{B}_{T}$ extends to some $D(\bar{c}_{1})$ which is t-isolating for $\bar{c}$ and $\bar{c}_{1}\subseteq\langle\bar{c}\rangle$ in every member of $D(\bar{c}_{1})]$.
@@
Proof. Necessity. Assuming that $G_{0}$ has a copy, say $\langle\bar{c}_{0}\rangle$, which is t-isolated by some $D_{0}(\bar{c}^{\prime}_{0})$ with $\bar{c}^{\prime}_{0}\subseteq\langle\bar{c}_{0}\rangle$, find an extension $D(\bar{c})\supseteq D_{0}(\bar{c}^{\prime}_{0})$ as in the formulation of WAP. Extend it to some t-isolating $D^{\prime}(\bar{c}^{\prime})$ for $\bar{c}$ with $\bar{c}^{\prime}\subseteq\langle\bar{c}\rangle$. It is easy to see that $D^{\prime}(\bar{c}^{\prime})$ witnesses WAP for $D_{0}(\bar{c}^{\prime}_{0})$. Let $G_{1}$ be the corresponding group $\langle\bar{c}\rangle$. Having $G_{2}=\langle\bar{c}_{2}\rangle$ and $G_{3}=\langle\bar{c}_{3}\rangle$ which contain $G_{1}$ and are t-isolated by some $D^{\prime}_{2}(\bar{c}^{\prime}_{2})$ and $D^{\prime}_{3}(\bar{c}^{\prime}_{3})$ with $\bar{c}^{\prime}_{2}\subseteq\langle\bar{c}_{2}\rangle$ and $\bar{c}^{\prime}_{3}\subseteq\langle\bar{c}_{3}\rangle$ (and where we may assume $\bar{c}\subseteq\bar{c}_{2}\cap\bar{c}_{3}$), we apply WAP for amalgamation $D^{\prime}_{2}(\bar{c}^{\prime}_{2})$ and $D^{\prime}_{3}(\bar{c}^{\prime}_{3})$ over $D_{0}(\bar{c}^{\prime}_{0})$. Assume that $D_{4}(\bar{c}_{4})$ is the corresponding amalgam. Then $\langle\bar{c}_{4}\rangle$ is a group containing $\bar{c}_{2}\cup\bar{c}_{3}$. By t-isolation, $\bar{c}_{2}$ and $\bar{c}_{3}$ generate subgroups of $\langle\bar{c}_{4}\rangle$ which are copies of $G_{2}$ and $G_{3}$.
Sufficiency. Let $D(\bar{c})\in\mathcal{B}_{T}$. Extend it to a t-isolating $D_{0}(\bar{c}^{\prime})$ for $\bar{c}$ with $\bar{c}^{\prime}\subseteq\langle\bar{c}\rangle$. Let $G_{0}=\langle\bar{c}\rangle$ be the corresponding t-isolated group. Find $G_{1}=\langle\bar{c}_{1}\rangle$ as in the formulation and the corresponding t-isolating diagram $D_{1}(\bar{c}^{\prime}_{1})$ with $\bar{c}^{\prime}_{1}\subseteq\langle\bar{c}_{1}\rangle$. We may assume that $\bar{c}^{\prime}\subseteq\bar{c}_{1}$. In order to verify amalgamation of extensions of $D_{1}(\bar{c}^{\prime}_{1})$ we first note that by density of t-isolating diagrams, we may restrict ourselves only to extensions that are t-isolating for some subtuples of their domains. If $D_{2}(\bar{c}^{\prime}_{2})$ and $D_{3}(\bar{c}^{\prime}_{3})$ are such extensions and $G_{2}=\langle\bar{c}_{2}\rangle$ and $G_{3}=\langle\bar{c}_{3}\rangle$ are the corresponding groups with $\bar{c}^{\prime}_{2}\subseteq\langle\bar{c}_{2}\rangle$ and $\bar{c}^{\prime}_{3}\subseteq\langle\bar{c}_{3}\rangle$, then by t-isolation, $G_{1}$ is a subgroup of both $G_{2}$ and $G_{3}$. Find $G_{4}$ which amalgamates $G_{2}$ and $G_{3}$ over $G_{0}$. It obviously belongs to some $D_{4}(\bar{c}_{4})]$ where $D_{4}(\bar{c}_{4})$ amalgamates $D_{2}(\bar{c}_{2})$ and $D_{3}(\bar{c}_{3})$ over $D_{0}(\bar{c}^{\prime})$. $\Box$
@@
It is worth noting that a slightly weaker form of the density assumption follows from WAP. Indeed, assume that there is $D(\bar{c})\in\mathcal{B}_{T}$ such that for every $D^{\prime}(\bar{c}^{\prime})\in\mathcal{B}_{T}$ extending $D(\bar{c})$ the diagram $D^{\prime}(\bar{c}^{\prime})$ is not t-isolating for $\bar{c}$ (where $\bar{c}^{\prime}\subseteq\langle\bar{c}\rangle$ is not assumed). Therefore, for such $D^{\prime}(\bar{c}^{\prime})$ two distinct groups can be realized as $\langle\bar{c}\rangle$ in some members of $D^{\prime}(\bar{c}^{\prime})]$. Thus $\mathcal{B}_{T}$ does not have WAP.
@@
As we already know, every finite group which extends to an enumerated group of $T$, is t-isolated (see Example 2.2). Using this we can conclude that if $T$ defines a variety of groups and all finitely generated groups satisfying $T$ are residually finite, then t-isolated diagrams are dense in $\mathcal{B}_{T}$. Indeed, assuming that $D(\bar{c})\in\mathcal{B}_{T}$ take a finite group $F$ generated by $\bar{c}$ and satisfying all relations which appear in $D(\bar{c})$. Let $\bar{c}_{1}$ be an enumeration of $F$. Then the multiplication table of $F$ (together with inequalities between distinct elements) is viewed as the corresponding diagram $D(\bar{c}_{1})$, where we also have $\bar{c}_{1}\subseteq\langle\bar{c}\rangle$. The diagram $D(\bar{c}_{1})$ works as an appropriate t-isolating extension of $D(\bar{c})$.
We also add here that under the assumptions of Remark 2.7, t-isolated groups satisfying $T$ are exactly finite ones. Indeed, if $G=\langle\bar{c}\rangle$ is infinite, then any basic set $D^{\prime}(\bar{c}^{\prime})]$ containing it, also contains a finite group.
@@ 2.2 Abelian groups
Abelian groups form a closed subspace of $\mathcal{G}$. We denote it by $\mathcal{G}_{Abel}$ where $Abel$ denotes the corresponding axiomatization. Under the scenario of case I of Section 1.2 we have $\mathcal{G}_{Abel}=\mathcal{X}_{M,Abel}$ .
The following theorem is proved in [6] and [7]. We give an easy and independent proof of it.
@@
Firstly, note that since finitely generated abelian groups are residually finite, Remark 2.7 implies that the t-isolated diagrams are dense in $\mathcal{B}_{Abel}$.
Let us verify the property that every t-isolated $G_{0}$ can be extended to a t-isolated $G_{1}$ such that any two t-isolated extensions of $G_{1}$ can be amalgamated over $G_{0}$. As we already mentioned after Remark 2.7, t-isolated abelian groups are exactly finite ones. Now the key point is the fact that the variety $\mathbb{G}_{Abel}$ has the amalgamation property. Indeed, if $A_{0}$ is a subgroup of $A_{1}$ and $\phi$ is an embedding of $A_{0}$ into $A_{2}$, then the group $(A_{1}\oplus A_{2})/\{(x-\phi(x))|x\in A_{0}\}$ is the corresponding amalgamation. $\Box$
2.3 Nilpotent groups of prime odd exponent
Let $c,p\in\omega\setminus\{0,1\}$ and $p$ be prime $\mathsf{max}(2,c)$. Nilpotent groups of degree $c$ and of exponent $p$ form a closed subspace of $\mathcal{G}$. We denote it by $\mathcal{G}_{c_{nil}p}$.
@@ 2.3 Nilpotent groups of prime odd exponent
Proof. Consider $\mathcal{G}_{c_{nil}p}$ as the space $\mathcal{X}_{M,c_{nil}p}$ under the scenario of case I of Section 1.2. According to Theorem 1.2 of [9] we have to verify that JEP and WAP hold for $\mathcal{B}_{c_{nil}p}$. Since nilpotent groups of degree $c$ and of exponent $p$ form a variety (we call it $\mathbb{G}_{c_{nil}p}$), JEP is satisfied: apply $\oplus$ for domains of the corresponding diagrams. In order to verify WAP we will apply Proposition 2.5.
First, note that since finitely generated nilpotent groups are residually finite (see Section 5.2 in [4]), the t-isolated diagrams are dense in $\mathcal{B}_{c_{nil}p}$. Furthermore, t-isolated groups of the variety $\mathbb{G}_{c_{nil}p}$ are exactly finite ones, see Remark 2.7 and the comment after it. Thus, Proposition 2.5 can be applied to finite groups. However, we need some additional algebraic material.
@@
Proof. Given $k\leq c$, we show how to build $G_{1}$ that satisfies the conclusion of the lemma for all $i\leq k$. To satisfy the conclusion of the lemma for $k=1$, take $G_{1}=G_{0}$. We now apply induction. Let $k<c$ and assume that we already have some $H\geq G_{0}$ satisfying the statement for all $i\leq k$. Let $G_{1}$ be a finite extension of $H$ with maximal $\Gamma_{k+1}(G_{1})\cap G_{0}$. Since for all $i\leq k$ we have $\Gamma_{i}(G_{1})\cap G_{0}=\Gamma_{i}(H)\cap G_{0}$, the statement of the conclusion of the lemma holds for $G_{1}$ and all $i\leq k+1$. $\Box$
In [2] A. Baudisch defines expansions of groups from $\mathbb{G}_{c_{nil}p}$ by additional unary predicates $P_{2},\ldots,P_{c}$. In order to present these expansions let $G$ be a group in $\mathbb{G}_{c_{nil}p}$ and let $Q_{1}(G)=P_{1}(G)=G$. When $1<n\leq c+1$ let $Q_{n}(G)=\langle\bigcup_{\ell+k=n}[P_{\ell}(G),P_{k}(G)]\rangle$ and $P_{n}(G)$ be a subgroup of $G$ with $Q_{n}(G)\subseteq P_{n}(G)\subseteq Z_{c+1-n}(G)$ and $P_{n}(G)\subseteq P_{n-1}(G)$. Let $\mathbb{G}^{P}_{c,p}$ be the class of all groups in $\mathbb{G}_{c_{nil}p}$ with additional predicates $P_{2}(G),\ldots,P_{c}(G)$ as described above. Let $L^{+}$ be the corresponding language that extends the language of groups. Note that if we extend a group $G\in\mathbb{G}_{c_{nil}p}$ by predicates $Q_{i}(G)=P_{i}(G)=\Gamma_{i}(G)$ for $1<i\leq c$ then we get an $L^{+}$-group from $\mathbb{G}^{P}_{c,p}$. The following statement is the main result of [2] (see Theorem 1.2 and the discussion after it).
The class of finite structures from $\mathbb{G}^{P}_{c,p}$ has the hereditary property and the amalgamation property.
@@
Consider $G_{1}$, $G_{2}$ and $G_{3}$ as $L^{+}$-structures defining $P_{i}$ on them by $\Gamma_{i}(G_{1})$, $\Gamma_{i}(G_{2})$ and $\Gamma_{i}(G_{3})$ respectively ($1<i\leq c$). Now for any $i\leq c$ let $P_{i}(G_{0})=\Gamma_{i}(G_{1})\cap G_{0}$. By the hereditary property $(G_{0},P_{2}(G_{0}),\ldots,P_{c}(G_{0}))$ is also an $L^{+}$-structure from $\mathbb{G}^{P}_{c,p}$.
Since $G_{1}$ is a subgroup of $G_{2}$ and $G_{3}$ then by Lemma 2.10, for any $i\leq c$, we have that $P_{i}(G_{0})=P_{i}(G_{1})\cap G_{0}=P_{i}(G_{2})\cap G_{0}=P_{i}(G_{3})\cap G_{0}$. In particular $(G_{0},P_{2}(G_{0}),\ldots,P_{c}(G_{0}))$ is an $L^{+}$-substructure of both $(G_{2},P_{2}(G_{2}),\ldots,P_{c}(G_{2}))$ and $(G_{3},P_{2}(G_{3}),\ldots,P_{c}(G_{3}))$. By Theorem 1.2 of [2] there is $L^{+}$-structure $G_{4}$ amalgamating $G_{2}$ and $G_{3}$ over $G_{0}$. Thus, the condition of Proposition 2.5 is satisfied. $\Box$